# Assessing treatment effects in observational data with missing confounders: A comparative study of practical doubly-robust and traditional missing data methods

### v1

@@ 3.1.6 Estimators
Census Benchmark BNMRK-C Ideal data $(Y,X,Z,W)$ All
Oracle Benchmark BNMRK-O Ideal data $(Y,X,Z,W)$ All
Complete-case CC $(RY,RX,RZ,R,RW)$ cOR
@@
3.1.7 Performance metrics
3.2 Simulation results
3.2.1 Base case (12% outcome probability, 40% MAR missing)
Figure 1 (Figure 2) shows the estimator relative performance for the cOR census (oracle) estimand for the base case simulations (detailed results in Tables S6S13). In the simple outcome and MAR scenario, the oracle and census estimands are the same and all methods, except the CFND and CC approaches, maintained the 95% nominal coverage. The GR estimator is consistent for the census estimand in all MAR cases because the analytic working model is the census model and so, as expected, maintained correct coverage for this parameter for these settings; unexpectedly, GR also maintained close to the correct coverage for both census and oracle estimands when the analytic working models both were misspecified (i.e. under the complex outcome and MAR scenario). GR also maintained the best efficiency for the census estimand and generally within about 5% of the best efficiency for the oracle estimand (i.e. lowest rRMSE) relative to the other methods for all 4 combinations of simple versus complex data generating scenarios. Conversely, MICE suffered from bias and inefficiency when the true outcome model was complex; this bias was worse for the census estimand. The MI procedures, even in the complex outcome scenario, were still more efficient than CC, IPW and the TMLE estimators; however, which imputation procedure did the best varied by setting. Generally the MI estimates were close when the outcome DGM was simple, with an apparent very slight advantage to MICE. When the outcome scenario was complex MI-RF had lower rRMSE than MICE for both the census and oracle estimands; MI-XGB tended to have lower rRMSE for the oracle cOR estimand while MI-RF had lower rRMSE for the census estimand (results for MI-XGB are only displayed in Tables S6S8, S10S12). The TMLE estimators were generally more efficient than IPW and CC, but not as efficient as MI and GR. Further, TMLE estimators, which were implemented as weighted two-phase estimators, suffered from some bias when the missing data model was more complex, with oracle coverage only around 92%; however, IPW fared much worse for the complex MAR model, with oracle (and nominal) coverage dropping to about 70%. TMLE-M was less efficient than TMLE-MTO for the complex outcome scenario. The CNFD method demonstrated severe bias, with nominal coverage under 20%, across all scenarios.
Figure 1: Synthetic Data MAR Simulation: Census clogOR. Comparing estimators of the census clogOR estimand for the base case with 40% missingness and 12% outcome probability, where the data generating models were simple or complex. Top graph: Percent Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at .$95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$. True clogOR values are 0.405 and 0.371 for simple and complex outcome models, respectively.
Figure 2: Synthetic Data MAR Simulation: Oracle clogOR. Comparing estimators of the oracle clogOR estimand for the base case with 40% missingness and 12% outcome probability, where the data generating models were simple or complex. Top graph: Percent Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at .$95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$. The true clogOR value is 0.405 for both simple and complex models.
3.2.2 Scenarios with varying outcome and missing data probabilities
3.2.3 MNAR Scenarios
Figure 3: Synthetic Data MNAR Simulation: Census clogOR. Comparing estimators of the census estimand with 40% confounder missingness and 12% outcome probability. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at .$95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$. True clogOR values are 0.405 for simple outcome models and 0.404 and 0.371 for complex outcome models for MNAR unobserved and MNAR value, respectively.
Figure 4: Synthetic Data MNAR Simulation: Oracle clogOR. Comparing estimators of the oracle estimand with 40% missingness and 12% outcome probability. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at .$95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$. The correct clogOR value is 0.405 across all scenarios.
4.1 Study cohort
4.2 Plasmode simulation methods
We seek to compare the same estimation methods described in Section 3.1.6 for four sets of plasmode simulation studies, where we vary the rarity of the outcome and the complexity of the underlying DGMs. We used two different DGMs that were derived by fitting different models to the KPWA cohort: 1) a parametric generalized linear model (GLM) (i.e., logistic regression) with a few interactions and 2) a tree-based approach allowing for complex interactions. These two DGMs allow us to investigate the performance of estimators under a simple and more complex DGM, where the latter was chosen to be have a functional form notably different than the typical linear parametric working model. In both DGMs, we used the same approach (i.e., GLM or trees) to estimate the propensity score, outcome regression, and missing-data models. We only consider the BNMRK-O estimator for the GLM DGM because there is no single oracle cOR of interest in the tree-based DGM.
We began by estimating the missing-data model using the entire sample of 112,770 individuals; in this model, the outcome was an indicator of whether the PHQ was observed or not. Next, we estimated the treatment propensity score and outcome regression models using the complete-data cohort (N=50,337). Importantly, using this complete-data cohort is convenient for plasmode simulation, but precludes making general conclusions about the larger population. Table S23 and Figures S36S38 describe the final parametric and tree-based models, respectively. With these models fixed, we generate plasmode data as follows. We use bootstrap sampling with replacement from the complete-data cohort (i.e., individuals with fully observed PHQ-9) to generate complete covariate data for each person for 1000 plasmode datasets of size N=50,337, thereby maintaining the complex data structure in the KPWA cohort. For each DGM approach, we use the fitted models to generate treatment given the plasmode sample of covariates, the outcome given the treatment and covariates, and then the missing data indicators for the PHQ-9 given the treatment, covariates and outcome for the 1000 plasmode datasets. The plasmode sampling for the simulations drew only from individuals with PHQ-9 measured (i.e., the complete-data cohort of size 50,337) and we use the generated missing data indicators to “hide” PHQ information for selected individuals.
Similar performance metrics as for the synthetic data simulations were calculated for each estimator across the plasmode simulations. Because the value of the DGM treatment effect estimands derived from the KPWA were close to null, bias instead of percent bias was calculated. The estimators returned results in $>98$% of plasmode datasets (Section S3.4).
4.3 Plasmode simulation results
Figure 5: Plasmode Data Simulation: Census mRD. Estimator relative performance for the census mRD estimand, across GLM and tree-based data generating models for the rare and common outcomes. Top graph: Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at .$95\pm 1.96\sqrt{\frac{.05\cdot.95}{1000}}.$. True values of mRD from lightest to darkest scenarios are 7.6E-4, 1.2E-4, -1.6E-2, and -6.2E-3.
Figure 6: Plasmode Data Simulation: Oracle mRD. Estimator relative performance for the oracle mRD estimand, across GLM and tree-based data generating models for the rare and common outcomes. Top graph: Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at .$95\pm 1.96\sqrt{\frac{.05\cdot.95}{1000}}.$. True values of mRD from lightest to darkest scenarios are 7.5E-4, -2.4E-4, -1.7E-2, and -7.6E-3.
5 Discussion
Our numerical experiments highlight the importance of first choosing a target estimand and then determining an estimation procedure. As expected, in cases where the working outcome regression model coincided with the data-generating outcome model and a simple MAR missing data model (the “base case” simulations), all methods performed well, with bias near zero and confidence interval coverage near the nominal 95% level. However, even in this standard case we noted several important findings. The doubly-robust GR approach, which reduces to a tailored re-weighting of complete cases, maintained nearly the same efficiency as MI, the method generally expected to be the most efficient for correctly-specified parametric models. While MI can be highly efficient, the MI method (e.g., MICE or MI-RF) that performs best can vary across scenarios, but we do not know the DGM scenario in practice; for this reason, GR is often a safer choice than MI. Perhaps less surprising, both the CFND and CC approaches were biased and highly inefficient; and IPW and TMLE were inefficient. Importantly, the CFND model — resulting from omitting the confounders with missing values — is always biased and is not recommended; all missing data approaches that we evaluated outperformed the CFND approach.
@@ 5 Discussion
Missing Scenario 2.5 MNAR with missingness dependent upon $U$ $\operatorname{logit}P(R=1\mid X=x,Z=z,Y=y,U=u)=\alpha_{0}+\alpha_{X}x+\alpha_{Z}^{\top}z+\alpha_{Y}y+\alpha_{U}^{\top}u$, $\alpha_{0}=-0.97$, $\alpha_{X}=\ln(2.5)$, $\alpha_{Z}=[\ln(1.5),\ln(1.5)]$, $\alpha_{Y}=\ln(2.5)$, [$\alpha_{Us},\alpha_{Uw}]=[\ln(2.5),\ln(1)]$
@@
Missing Scenario 2.7 Same MNAR as 2.5 with 80% missing $\operatorname{logit}P(R=1\mid X=x,Z=z,Y=y,U=u)=\alpha_{0}+\alpha_{X}x+\alpha_{Z}^{\top}z+\alpha_{Y}y+\alpha_{U}^{\top}u$, $\alpha_{X}=\ln(2.5)$, $\alpha_{Z}=[\ln(1.5),\ln(1.5)]$, $\alpha_{Y}=\ln(2.5)$, [$\alpha_{Us},\alpha_{Uw}]=[\ln(2.5),\ln(1)]$, $\alpha_{0}=1.28$
@@
Supplemental Figure S4: Synthetic Data MAR Simulation: Census mRD. Comparing estimators of the census estimand with 40% confounder missingness MAR and 12% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at .$95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$. True mRD values are 0.040 and 0.037 for simple and complex outcome models, respectively.
Supplemental Figure S5: Synthetic Data MAR Simulation: oracle mRD. Comparing estimators of the oracle truth with 40% missingness MAR and 12% incidence. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at .$95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$. True mRD values are 0.040 and 0.031 for simple and complex outcome models, respectively.
@@
Supplemental Figure S6: Synthetic Data MAR Simulation: Census mlogRR. Comparing estimators of the census estimand with 40% confounder missingness and 12% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at .$95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$. True values of mlogRR are 0.334 and 0.315 for simple and complex outcome models, respectively.
Supplemental Figure S7: Synthetic Data MAR Simulation: Oracle mlogRR. Comparing estimators of the oracle estimand with 40% confounder missingness and 12% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at .$95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$. True values of mlogRR are 0.334 and 0.271 for simple and complex outcome models, respectively.
@@
Supplemental Figure S8: Synthetic Data MAR Simulation: Census clogOR. Comparing estimators of the census estimand with 80% missingness and 12% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at .$95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$. True clogOR values are 0.405 and 0.371 for simple and complex outcome models, respectively.
Supplemental Figure S9: Synthetic Data MAR Simulation: Oracle clogOR. Comparing estimators of the oracle estimand with 80% confounder missingness MAR and 12% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at .$95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$. The true clogOR value is 0.405 for both simple and complex outcome models.
@@
Supplemental Figure S10: Synthetic Data MAR Simulation: Census mRD. Comparing estimators of the census truth with 80% confounder missingness and 12% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at .$95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$. True mRD values are 0.040 and 0.037 for simple and complex outcome models, respectively.
Supplemental Figure S11: Synthetic Data MAR Simulation: Oracle mRD. Comparing estimators of the oracle estimand with 80% confounder missingness and 12% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at .$95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$. True mRD values are 0.040 and 0.031 for simple and complex outcome models, respectively.
@@
Supplemental Figure S12: Synthetic Data MAR Simulation: Census mlogRR. Comparing estimators of the census truth with 80% confounder missingness and 12% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at .$95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$. True values of mlogRR are 0.334 and 0.315 for simple and complex outcome models, respectively.
Supplemental Figure S13: Synthetic Data MAR Simulation: Oracle mlogRR. Comparing estimators of the oracle estimand with 80% confounder missingness and 12% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at .$95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$. True values of mlogRR are 0.334 and 0.271 for simple and complex outcome models, respectively.
@@
Supplemental Figure S14: Synthetic Data MAR Simulation: Census clogOR. Comparing estimators of the census estimand with 80% confounder missingness MAR and 5% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at .$95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$. True clogOR values are 0.405 and 0.380 for simple and complex outcome models, respectively.
Supplemental Figure S15: Synthetic Data MAR Simulation: Oracle clogOR. Comparing estimators of the oracle estimand with 80% confounder missingness MAR and 5% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at .$95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$. The true clogOR value is 0.405 for both simple and complex outcome models.
@@
Supplemental Figure S16: Synthetic Data MAR Simulation: Census mRD. Comparing estimators of the census estimand with 80% confounder missingness and 5% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at .$95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$. The true mRD value is 0.019 for both simple and complex outcome models.
Supplemental Figure S17: Synthetic Data MAR Simulation: Oracle mRD. Comparing estimators of the oracle estimand with 80% confounder missingness and 5% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at .$95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$. True mRD values are 0.019 and 0.015 for simple and complex outcome models, respectively.
@@
Supplemental Figure S18: Synthetic Data MAR Simulation: Census mlogRR. Comparing estimators of the census estimand with 80% confounder missingness and 5% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at .$95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$. True values of mlogRR are 0.370 and 0.349 for simple and complex outcome models, respectively.
Supplemental Figure S19: Synthetic Data MAR Simulation: Oracle mlogRR. Comparing estimators of the oracle estimand with 80% confounder missingness and 5% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at .$95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$. True values of mlogRR are 0.370 and 0.281 for simple and complex outcome models, respectively.
@@
Supplemental Figure S20: Synthetic Data MNAR Simulation: Census mRD. Comparing estimators of the census estimand with 40% confounder missingness and 12% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at .$95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$. The true mRD values are 0.040 for simple outcome models, and 0.038 and 0.037 for complex outcome models for MNAR unobserved and MNAR value, respectively.
Supplemental Figure S21: Synthetic Data MNAR Simulation: Oracle mRD. Comparing estimators of the oracle truth with 40% confounder missingness and 12% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at .$95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$. The true mRD values are 0.040 for simple outcome models, and 0.038 and 0.031 for complex outcome models for MNAR unobserved and MNAR value, respectively.
@@
Supplemental Figure S22: Synthetic Data MNAR Simulation: Census mlogRR. Comparing estimators of the census estimand with 40% confounder missingness and 12% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at .$95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$. True values of mlogRR are 0.334 for simple outcome models and 0.339 and 0.315 for complex outcome models for MNAR unobserved and MNAR value, respectively.
Supplemental Figure S23: Synthetic Data MNAR Simulation: Oracle mlogRR. Comparing estimators of the oracle estimand with 40% confounder missingness and 12% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at .$95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$. True values of mlogRR are 0.334 for simple outcome models and 0.339 and 0.271 for complex outcome models for MNAR unobserved and MNAR value, respectively.

### v2

@@ 5 Discussion
Census Benchmark BNMK-C Ideal data $(Y,X,Z,W)$ All
Oracle Benchmark BNMK-O Ideal data $(Y,X,Z,W)$ All
Complete-case $C$ $(RY,RX,RZ,R,RW)$ cOR
@@
Missing Scenario 2.5 MNAR with missingness dependent upon $U$ $\operatorname{logit}P(R=1\mid X=x,Z=z,Y=y,U=u)=\alpha_{0}+\alpha_{X}x+\alpha_{Z}^{\top}z+\alpha_{Y}y+\alpha_{U}^{\top}u$, $\alpha_{0}=-0.97$, $\alpha_{X}=\ln(2.5)$, $\alpha_{Z}=[\ln(1.5),\ln(1.5)]$, $\alpha_{Y}=\ln(2.5)$, $\alpha_{Us},\alpha_{Uw}]=[\ln(2.5),\ln(1)]$
@@
Missing Scenario 2.7 Same MNAR as 2.5 with 80% missing $\operatorname{logit}P(R=1\mid X=x,Z=z,Y=y,U=u)=\alpha_{0}+\alpha_{X}x+\alpha_{Z}^{\top}z+\alpha_{Y}y+\alpha_{U}^{\top}u$, $\alpha_{X}=\ln(2.5)$, $\alpha_{Z}=[\ln(1.5),\ln(1.5)]$, $\alpha_{Y}=\ln(2.5)$, $\alpha_{Us},\alpha_{Uw}]=[\ln(2.5),\ln(1)]$, $\alpha_{0}=1.28$
@@
Supplemental Figure S4: Synthetic Data MAR Simulation: Census mRD. Comparing estimators of the census estimand with 40% confounder missingness MAR and 12% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at $95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$ True mRD values are 0.040 and 0.037 for simple and complex outcome models, respectively.
Supplemental Figure S5: Synthetic Data MAR Simulation: oracle mRD. Comparing estimators of the oracle truth with 40% missingness MAR and 12% incidence. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at $95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$ True mRD values are 0.040 and 0.031 for simple and complex outcome models, respectively.
@@
Supplemental Figure S6: Synthetic Data MAR Simulation: Census mlogRR. Comparing estimators of the census estimand with 40% confounder missingness and 12% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at $95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$ True values of mlogRR are 0.334 and 0.315 for simple and complex outcome models, respectively.
Supplemental Figure S7: Synthetic Data MAR Simulation: Oracle mlogRR. Comparing estimators of the oracle estimand with 40% confounder missingness and 12% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at $95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$ True values of mlogRR are 0.334 and 0.271 for simple and complex outcome models, respectively.
@@
Supplemental Figure S8: Synthetic Data MAR Simulation: Census clogOR. Comparing estimators of the census estimand with 80% missingness and 12% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at $95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$ True clogOR values are 0.405 and 0.371 for simple and complex outcome models, respectively.
Supplemental Figure S9: Synthetic Data MAR Simulation: Oracle clogOR. Comparing estimators of the oracle estimand with 80% confounder missingness MAR and 12% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at $95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$ The true clogOR value is 0.405 for both simple and complex outcome models.
@@
Supplemental Figure S10: Synthetic Data MAR Simulation: Census mRD. Comparing estimators of the census truth with 80% confounder missingness and 12% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at $95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$ True mRD values are 0.040 and 0.037 for simple and complex outcome models, respectively.
Supplemental Figure S11: Synthetic Data MAR Simulation: Oracle mRD. Comparing estimators of the oracle estimand with 80% confounder missingness and 12% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at $95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$ True mRD values are 0.040 and 0.031 for simple and complex outcome models, respectively.
@@
Supplemental Figure S12: Synthetic Data MAR Simulation: Census mlogRR. Comparing estimators of the census truth with 80% confounder missingness and 12% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at $95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$ True values of mlogRR are 0.334 and 0.315 for simple and complex outcome models, respectively.
Supplemental Figure S13: Synthetic Data MAR Simulation: Oracle mlogRR. Comparing estimators of the oracle estimand with 80% confounder missingness and 12% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at $95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$ True values of mlogRR are 0.334 and 0.271 for simple and complex outcome models, respectively.
@@
Supplemental Figure S14: Synthetic Data MAR Simulation: Census clogOR. Comparing estimators of the census estimand with 80% confounder missingness MAR and 5% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at $95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$ True clogOR values are 0.405 and 0.380 for simple and complex outcome models, respectively.
Supplemental Figure S15: Synthetic Data MAR Simulation: Oracle clogOR. Comparing estimators of the oracle estimand with 80% confounder missingness MAR and 5% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at $95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$ The true clogOR value is 0.405 for both simple and complex outcome models.
@@
Supplemental Figure S16: Synthetic Data MAR Simulation: Census mRD. Comparing estimators of the census estimand with 80% confounder missingness and 5% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at $95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$ The true mRD value is 0.019 for both simple and complex outcome models.
Supplemental Figure S17: Synthetic Data MAR Simulation: Oracle mRD. Comparing estimators of the oracle estimand with 80% confounder missingness and 5% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at $95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$ True mRD values are 0.019 and 0.015 for simple and complex outcome models, respectively.
@@
Supplemental Figure S18: Synthetic Data MAR Simulation: Census mlogRR. Comparing estimators of the census estimand with 80% confounder missingness and 5% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at $95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$ True values of mlogRR are 0.370 and 0.349 for simple and complex outcome models, respectively.
Supplemental Figure S19: Synthetic Data MAR Simulation: Oracle mlogRR. Comparing estimators of the oracle estimand with 80% confounder missingness and 5% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at $95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$ True values of mlogRR are 0.370 and 0.281 for simple and complex outcome models, respectively.
@@
Supplemental Figure S20: Synthetic Data MNAR Simulation: Census mRD. Comparing estimators of the census estimand with 40% confounder missingness and 12% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at $95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$ The true mRD values are 0.040 for simple outcome models, and 0.038 and 0.037 for complex outcome models for MNAR unobserved and MNAR value, respectively.
Supplemental Figure S21: Synthetic Data MNAR Simulation: Oracle mRD. Comparing estimators of the oracle truth with 40% confounder missingness and 12% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at $95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$ The true mRD values are 0.040 for simple outcome models, and 0.038 and 0.031 for complex outcome models for MNAR unobserved and MNAR value, respectively.
@@
Supplemental Figure S22: Synthetic Data MNAR Simulation: Census mlogRR. Comparing estimators of the census estimand with 40% confounder missingness and 12% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at $95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$ True values of mlogRR are 0.334 for simple outcome models and 0.339 and 0.315 for complex outcome models for MNAR unobserved and MNAR value, respectively.
Supplemental Figure S23: Synthetic Data MNAR Simulation: Oracle mlogRR. Comparing estimators of the oracle estimand with 40% confounder missingness and 12% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at $95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$ True values of mlogRR are 0.334 for simple outcome models and 0.339 and 0.271 for complex outcome models for MNAR unobserved and MNAR value, respectively.